LjTheBaller16:

Nicole wants to buy a pair of jeans. The original cost of the jeans is $42.50, and the markup is 10 percent. How much will she have to pay for the jeans?

XxAngelGirlxX:

42.50 * 10% or 0.10 = 4.25

foolghost14:

lol

XioGonz:

42.50 / 10% = 425 =$4.25.\ 42.50 + 4.25 = 46.75 Your final answer will not be 4.25 Your final answer is 46.75